If 10^x/2 + 10^-x/2 =4, prove that x = log(7 + 4√3)


1
Expert's answer
2021-03-09T07:50:21-0500

Let "10^{x}=u, u>0." Then


"u+u^{-1}=14"

"u^2-14u+1=0"

"u=7\pm4\sqrt{3}"

"4\sqrt{3}=\sqrt{48}<7"


"u_1=7-4\sqrt{3},u_2=7+4\sqrt{3}"


"x_1=\log{(7-4\sqrt{3})}"


"x_2=\log{(7+4\sqrt{3})}"